# Language-based Valence and Arousal Expressions between the United States and China: a Cross-Cultural Examination

### v4

@@ Abstract
While affective expressions on social media have been extensively studied, most research has focused on the Western context. This paper explores cultural differences in affective expressions by comparing valence and arousal on Twitter/X (geolocated to the US) and Sina Weibo (in Mainland China). Using the NRC-VAD lexicon to measure valence and arousal, we identify distinct patterns of emotional expression across both platforms. Our analysis reveals a functional representation between valence and arousal, showing a negative offset in contrast to traditional lab-based findings which suggest a positive offset. Furthermore, we uncover significant cross-cultural differences in arousal, with US users displaying higher emotional intensity than Chinese users, regardless of the valence of the content. Finally, we conduct a comprehensive language analysis correlating n-grams and LDA topics with affective dimensions to deepen our understanding of how language and culture shape emotional expression. These findings contribute to a more nuanced understanding of affective communication across cultural and linguistic contexts on social media.
@@ 1 Introduction
Subjective expressions of affect (how we feel) play a crucial role in understanding learning outcomes in individuals Hourihan et al. (2017), their perceptions Gorn et al. (2001), well-being Xu et al. (2015), and mental and physical health Cohen and Pressman (2006). Multiple theoretical and empirical works have, therefore, examined the underlying dimensions of affect and their relationships. While there are several models of affective structure, Russell’s two-dimensional circumplex model is the most widely recognized, where orthogonal valence (pleasant to unpleasant) and arousal (high to low activation) are represented as the horizontal and vertical axes Russell (1980); Yik et al. (1999).
@@
To address these limitations, researchers advocate for studies that go beyond self-reports and focus on behaviors Baumeister et al. (2007), especially on social media. Social media data offer a naturalistic and ecological setting to capture individuals’ emotions and, despite potential social desirability bias, have been shown to reliably estimate well-being Jaidka et al. (2020); Liou et al. (2023), sentiment Preoţiuc-Pietro et al. (2016), and personality Schwartz et al. (2013); Havaldar et al. (2024).
This paper examines the cross-cultural difference on the affective expressions with functional relationship between valence and arousal by analyzing natural language expressions from Twitter/X (geolocated to the US) and Sina Weibo (in Mainland China) posts. The pipeline of our work is shown in Table 1. The study has three key contributions:
@@ 2.1 Data Preparation
Weibo, unlike Twitter, does not offer an API tool for obtaining random samples over time. So, starting with a random set of individuals from a public dataset Guntuku et al. (2019a), Weibo posts were gathered using a breadth-first search method on users. We obtained over 29 million posts from 2014 from 859,054 people on Weibo. Gender and age were collected from self-reported demographic information on their Weibo profile. Subsetting to users with more than 500 words and with a reasonable self-reported age (<100 years) and gender, the dataset consisted of 668,257 Weibo posts from 8,731 users. 500 words were found to be the minimum threshold to obtain reliable psychological estimates from individuals’ language Eichstaedt et al. (2021); Jaidka et al. (2018).
Based on the gender and age distribution of Weibo and Twitter users, we built propensity-score-based matched samples, resulting in 2,191 users each on both platforms with at least 500 words. These matched users had 2.4 million posts on Twitter and 177,042 posts on Weibo. In our matched dataset 67.1% self-reported as being female and 32.9% as male, and the mean age was 26.9 (s.d. 8.8). On Twitter, there were on average 15.6 (s.d. 2.8) words per user and Weibo had 57.3 (s.d. 15.4) words per user. The differences in word counts are driven by the lack of character limits to posts on Weibo.
@@ 2.2 Valence & Arousal Measurement
The circumplex and vector models of emotion have been broadly used for representing affective states Russell (1980); Bradley et al. (1992). In these two-dimensional models, valence is the x-axis, expressing pleasantness and unpleasantness, attractiveness and aversiveness, joy, and sorrow Frijda (1986). Arousal is the y-axis, describing the degree of wakefulness, boredom, excitement, and calm. These models allow any affective state, emotion, word, or expression to be represented as a point in the space, regardless of the difference in language, country, or culture.
@@ 2.3 Evaluation of Functional Relationship
Kuppens et al., 2013 showed six possible functional relationships between valence and arousal. These models are independence (Model 1), Linear Relation (Model 2), Symmetric V-Shaped Relation (Model 3), and Asymmetric V-Shaped Relations, including asymmetric interception (Model 4), asymmetric slope (Model 5), and asymmetric interception and slope (Model 6). The models’ functional representations are shown below:
@@
While applying the six models, we use mixed effects models to fit the datasets. We assume there is a fixed relationship between valence and arousal across all posts, while the average level of arousal may vary from user to user. The regression models can correctly represent the relationship between the two variables by setting within-person differences as the random effect. In model comparison, to cover both fixed and random effects, we use conditional $R^{2}$ to represent the proportion of variance explained by the entire model.
@@ 2.4 Social Media Language Analysis
We extract two open-vocabulary features from Twitter and Weibo: n-grams and topics. We extract contiguous sequences of one or two words (1-2 grams, Kern et al., 2014; Andrew Schwartz et al., 2013) with pointwise mutual information (PMI = 3; Church and Hanks, 1990). This resulted in unique unigrams and bigrams set of 10,477 for Weibo and 12,798 for Twitter. We extracted the normalized distribution of the n-grams for each post in the Weibo and Twitter datasets. We then used 2000 topics generated using Latent Dirichlet Allocation (LDA, Blei et al., 2003) as the second feature to represent users’ language in our Twitter and Weibo datasets. We utilized topics generated on much larger datasets to favor high diversity and coverage. 2,000 English topics generated a corpus of approximately 18 million Facebook updates with alpha set to 0.30 to favor fewer topics per document. These have been shown to perform well across multiple platforms Eichstaedt et al. (2015). 2,000 Mandarin topics were generated on 29 million Weibo posts with similar parameters set in Mallet. Inherently, each topic is realized as a set of words with probabilities. Every post is thus scored in terms of its probability of containing each of the 2000 topics, $p(topic,post)$, which is derived from their probability of containing a word, $p(word|post)$, and the probability of the words being in given topics, $p(topic|word)$.
@@ 3.2 Social Media Language Analysis
We further compare Twitter and Weibo’s DLA results for topics in Figure 4 and 5. Twitter users had relaxing weekend (’weekend’, ’awesome’, ’amazing’, ’great’, ’retreat’), celebration of events (’birthday’, ’wishes’, ’happy’, ’present’, ’wished’), luck and achievement (’win’, ’won’, ’contest’, ’prize’, ’lottery’) for positive valence high arousal. Conversely, Weibo users discussed affectionate bonding (’love’, ’hopeless’, ’willing’, ’protective’, ’friendly’, where hopeless means love in deep) to express their feelings, particularly in the context of festivals and celebrations (’new year’, ’red envelope’) and interests in celebrities and TV shows (’celebrities’,’singer’). For positive valence low arousal, Twitter users usually talked about relaxing routines (’day’, ’today’, ’good’, ’chilled’) and sleep (’night’, ’sleep’, ’tonight’, ’rest’, ’hoping’). Besides, Weibo users shared family reunion (’home’, ’return’, ’mother’, ’family’, ’new year’) and savory cuisines (’dish’, ’meat’, ’delicious’, ’soup’, ’dish’). When expressing strong negative feelings, Twitter users mainly used profanities (’fucking’, ’fuck’, ’shit’, ’pissed’, ’bullshit’) to convey intense emotions, while Weibo users discussed law enforcement and criminal investigation (’police’, ’crime’, ’suspect’, ’caught’, ’case’). Additionally, Weibo discussions on negative high arousal included the use of emojis (’sweat’) and negative emotions (’shocking’, ’hurt’, ’give up’). Concerning negative valence low arousal, Twitter users usually showed personal negative feelings like tiredness (’tired’, ’sleepy’, ’sleep’, ’sooo’, ’ugh’), engaged in discussions about daily activities (’hair’, ’cut’, ’short’, ’haircut’, ’cutting’) and mentioned words related to time (’hour’, ’minute’). Similarly, Weibo users also mentioned sleep (’sleep’, ’awake’, ’bed’).
@@ 4 Discussion
Similarly, past literature suggests that high-arousal emotions serve as an effective means of influencing others in the West Tsai (2007), while low-arousal emotions serve as an effective means of adjusting and conforming to others in the East Markus and Kitayama (1991). We found that low arousal emotions in Weibo were used to influence others through sharing wisdom and conclusions about life and love. On the high arousal-high positive affect sphere, Twitter users celebrated more personal events, while Weibo users talked more excitedly about celebrities and current events. Therefore, it is likely that while the ideal affect preference translates into affective expressions about personal experiences in the East, discussion of media culture is exempt from such norms: for instance, while it may be frowned upon to act too excited about personal events, the same restrictions are not in place when expressing excitement about celebrities and cultural events. As such, our findings provide a novel insight into our understanding of norm differences in affective expression in East vs West.
Similarly, looking at the difference in negativity bias for Twitter and Weibo, while Twitter users use profanity primarily, Weibo users tend to use words with much lower intensity, confirming the assumption that Chinese users try to avoid expressing extreme emotions.
@@ 6 Limitations
Translation: Lexica need to be adapted to the cultures to measure psychological phenomenon accurately. We tried using Chinese valence-arousal words (CVAW, Lee et al. (2022)). However, we did not proceed further as the methods of building NRC-VAD (for English) and CVAW (for Chinese) lexica were different and could cause misalignment. We wanted to control for such differences by choosing a lexicon that has sufficient coverage while also being used in multiple prior works across both languages Wenjia et al. (2023); Mohammad (2016); Li et al. (2019); Chen et al. (2019); Das and Dutta (2021). Further, with over 20K entries, NRC-VAD is the largest manually created emotion lexicon that has translations across several languages.
Censorship: We acknowledge that censorship on Sina Weibo is a challenge. Despite this issue, Weibo has been successfully used across multiple studies to understand different psychological outcomes (e.g. affect, stress, depression; Pan et al. (2021); Tang et al. (2019); Li et al. (2014))
@@ 7 Ethics
The data collection in our study were approved by the Institutional Review Board (IRB) of our institution.
@@
Appendix A Explanation of Conditional $R^{2}$
@@
The method we used in finding functional relationships between valence and arousal is a follow-up analysis using the same method used in Kuppens, Tuerlinckx, Russell, and Barrett (2013). Table 2 in the paper listed BIC, PostP, and the best $R^{2}$ for different studies and datasets, which are consistent with our result - relatively low $R^{2}$ scores. In our paper, we expand this work using social media data to see if a similar conclusion can hold across cultures with the functional relationship between valence and arousal.
@@
Appendix B Details of Valence-Arousal Relation Models
@@
Appendix C Removal of NRC-VAD Lexicons
@@
Appendix D Pre-Translated Weibo Figures

### v5

@@ Abstract
While affective expressions on social media have been extensively studied, most research has focused on the Western context. This paper explores cultural differences in affective expressions by comparing valence and arousal on Twitter/X (geolocated to the US) and Sina Weibo (in Mainland China). Using the NRC-VAD lexicon to measure valence and arousal, we identify distinct patterns of emotional expression across both platforms. Our analysis reveals a functional representation between valence and arousal, showing a negative offset in contrast to traditional lab-based findings which suggest a positive offset. Furthermore, we uncover significant cross-cultural differences in arousal, with US users displaying higher emotional intensity than Chinese users, regardless of the valence of the content. Finally, we conduct a comprehensive language analysis correlating n-grams and LDA topics with affective dimensions to deepen our understanding of how language and culture shape emotional expression. These findings contribute to a more nuanced understanding of affective communication across cultural and linguistic contexts on social media. [^1] For future research, we release our dataset at https://github.com/JeffreyCh0/X_Weibo_Valence_Arousal
@@ 1 Introduction
Subjective expressions of affect (how we feel) play a crucial role in understanding learning outcomes in individuals Hourihan et al. (2017), their perceptions Gorn et al. (2001), well-being Xu et al. (2015), and mental and physical health Cohen and Pressman (2006). Multiple theoretical and empirical works have, therefore, examined the underlying dimensions of affect and their relationships. While there are several models of affective structure, Russell’s two-dimensional circumplex model is the most widely recognized, where orthogonal valence (pleasant to unpleasant) and arousal (high to low activation) are represented as the horizontal and vertical axes [^2] See Russell’s two-dimensional circumplex model in Figure 6. Russell (1980); Yik et al. (1999).
@@
To address these limitations, researchers advocate for studies that go beyond self-reports and focus on behaviors Baumeister et al. (2007), especially on social media. Social media data offer a naturalistic and ecological setting to capture individuals’ emotions and, despite potential social desirability bias, have been shown to reliably estimate well-being Jaidka et al. (2020); Liou et al. (2023), sentiment Preoţiuc-Pietro et al. (2016), and personality Schwartz et al. (2013a); Havaldar et al. (2024).
This paper examines the cross-cultural difference on the affective expressions with functional relationship between valence and arousal by analyzing natural language expressions from Twitter/X (geolocated to the US) and Sina Weibo (in Mainland China) posts. The pipeline of our work is shown in Figure 1. The study has three key contributions:
@@ 2.1 Data Preparation
Weibo, unlike Twitter, does not offer an API tool for obtaining random samples over time. So, starting with a random set of individuals from a public dataset Guntuku et al. (2019a), Weibo posts were gathered using a breadth-first search method on users [^3] We started with a random set of users, and expanded to all their friends (bidirectional, similar to followers + following), and we repeated the process.. We obtained over 29 million posts from 2014 from 859,054 people on Weibo. Gender and age were collected from self-reported demographic information on their Weibo profile. Subsetting to users posted more than 500 words and with a reasonable self-reported age (<100 years) and gender, the dataset consisted of 668,257 Weibo posts from 8,731 users. 500 words were found to be the minimum threshold to obtain reliable psychological estimates from individuals’ language Eichstaedt et al. (2021); Jaidka et al. (2018).
Based on the gender and age distribution of Weibo and Twitter users, we built propensity-score-based matched samples, resulting in 2,191 users each on both platforms with at least 500 words [^4] We use propensity score matching for its nuanced handling of continuous variables and its allowance for quantitative assessment of covariate balance between matched groups. Rosenbaum and Rubin (1983). These matched users had 2.4 million posts on Twitter and 177,042 posts on Weibo. In our matched dataset 67.1% self-reported as being female and 32.9% as male, and the mean age was 26.9 (s.d. 8.8). On Twitter, there were on average 15.6 (s.d. 2.8) words per user and Weibo had 57.3 (s.d. 15.4) words per user. The differences in word counts are driven by the lack of character limits to posts on Weibo.
@@ 2.2 Valence & Arousal Measurement
The circumplex and vector models of emotion have been broadly used for representing affective states [^5] See Russell’s two-dimensional circumplex model in Figure 6. Russell (1980); Bradley et al. (1992). In these two-dimensional models, valence is the x-axis, expressing pleasantness and unpleasantness, attractiveness and aversiveness, joy, and sorrow Frijda (1986). Arousal is the y-axis, describing the degree of wakefulness, boredom, excitement, and calm. These models allow any affective state, emotion, word, or expression to be represented as a point in the space, regardless of the difference in language, country, or culture.
@@ 2.3 Evaluation of Functional Relationship
Kuppens et al., 2013 showed six possible functional relationships between valence and arousal. These models are independence (Model 1), Linear Relation (Model 2), Symmetric V-Shaped Relation (Model 3), and Asymmetric V-Shaped Relations, including asymmetric interception (Model 4), asymmetric slope (Model 5), and asymmetric interception and slope (Model 6). The models’ functional representations are shown below [^6] See full examples on our dataset on Figure 10.:
@@
While applying the six models, we use mixed effects models to fit the datasets. We assume there is a fixed relationship between valence and arousal across all posts, while the average level of arousal may vary from user to user. The regression models can correctly represent the relationship between the two variables by setting within-person differences as the random effect. In the model comparison, to cover both fixed and random effects, we use conditional $R^{2}$ to represent the proportion of variance explained by the entire model.
@@ 2.4 Social Media Language Analysis
We extract two open-vocabulary features from Twitter and Weibo: n-grams and topics. N-grams help capture common word patterns and phrase structures that reflect how emotions are expressed in everyday language, allowing us to identify culturally specific linguistic cues tied to affective dimensions. Meanwhile, Latent Dirichlet Allocation (LDA, Blei et al., 2003) uncovers underlying topics within the text, revealing thematic contexts that influence emotional expression. We choose LDA for topic modeling because it has better explainability and computational efficiency than other modern models like Top2Vec and BERTopic Angelov (2020); Grootendorst (2022). By analyzing both n-grams and topics, we can better understand the interplay between language, culture, and emotion, providing a richer, data-driven perspective on cross-cultural affective communication.
We collect contiguous sequences of one or two words (1-2 grams, Kern et al., 2014; Andrew Schwartz et al., 2013) with pointwise mutual information (PMI = 3; Church and Hanks, 1990). This resulted in unique unigrams and bigrams set of 10,477 for Weibo and 12,798 for Twitter. We extracted the normalized distribution of the n-grams for each post in the Weibo and Twitter datasets. We then used 2,000 topics generated using LDA as the second feature to represent users’ language in our Twitter and Weibo datasets Schwartz et al. (2013b). We utilized topics generated on much larger datasets to favor high diversity and coverage. 2,000 English topics generated a corpus of approximately 18 million Facebook updates with alpha set to 0.30 to favor fewer topics per document. These have been shown to perform well across multiple platforms Eichstaedt et al. (2015). 2,000 Mandarin topics were generated on 29 million Weibo posts with similar parameters set in Mallet McCallum (2002). Inherently, each topic is realized as a set of words with probabilities. Every post is thus scored in terms of its probability of containing each of the 2,000 topics, $p(topic,post)$, which is derived from their probability of containing a word, $p(word|post)$, and the probability of the words being in given topics, $p(topic|word)$.
@@ 3.2 Social Media Language Analysis
We further compare Twitter and Weibo’s LDA results for topics in Figure 4 and 5. Twitter users had relaxing weekend (’weekend’, ’awesome’, ’amazing’, ’great’, ’retreat’), celebration of events (’birthday’, ’wishes’, ’happy’, ’present’, ’wished’), luck and achievement (’win’, ’won’, ’contest’, ’prize’, ’lottery’) for positive valence high arousal. Conversely, Weibo users discussed affectionate bonding (’love’, ’hopeless’, ’willing’, ’protective’, ’friendly’, where hopeless means love in deep) to express their feelings, particularly in the context of festivals and celebrations (’new year’, ’red envelope’) and interests in celebrities and TV shows (’celebrities’,’singer’). For positive valence low arousal, Twitter users usually talked about relaxing routines (’day’, ’today’, ’good’, ’chilled’) and sleep (’night’, ’sleep’, ’tonight’, ’rest’, ’hoping’). Besides, Weibo users shared family reunion (’home’, ’return’, ’mother’, ’family’, ’new year’) and savory cuisines (’dish’, ’meat’, ’delicious’, ’soup’, ’dish’). When expressing strong negative feelings, Twitter users mainly used profanities (’fucking’, ’fuck’, ’shit’, ’pissed’, ’bullshit’) to convey intense emotions, while Weibo users discussed law enforcement and criminal investigation (’police’, ’crime’, ’suspect’, ’caught’, ’case’). Additionally, Weibo discussions on negative high arousal included the use of emojis (’sweat’) and negative emotions (’shocking’, ’hurt’, ’give up’). Concerning negative valence low arousal, Twitter users usually showed personal negative feelings like tiredness (’tired’, ’sleepy’, ’sleep’, ’sooo’, ’ugh’), engaged in discussions about daily activities (’hair’, ’cut’, ’short’, ’haircut’, ’cutting’) and mentioned words related to time (’hour’, ’minute’). Similarly, Weibo users also mentioned sleep (’sleep’, ’awake’, ’bed’).
@@ 4 Discussion
Similarly, past literature suggests that high-arousal emotions serve as an effective means of influencing others in the West Tsai (2007), while low-arousal emotions serve as an effective means of adjusting and conforming to others in the East Markus and Kitayama (1991). We found that low arousal emotions in Weibo were used to create a sense of comfort and connection through themes related to nature, family, daily life activities, and light-hearted entertainment. On the high arousal-high positive affect sphere, Twitter users celebrated more personal events, while Weibo users talked more excitedly about celebrities and current events. Therefore, it is likely that while the ideal affect preference translates into affective expressions about personal experiences in the East, discussion of media culture is exempt from such norms: for instance, while it may be frowned upon to act too excited about personal events, the same restrictions are not in place when expressing excitement about celebrities and cultural events. As such, our findings provide a novel insight into our understanding of norm differences in affective expression in East vs West.
Similarly, looking at the difference in negativity bias for Twitter and Weibo, while Twitter users use profanity primarily, Weibo users tend to use words with much lower intensity, confirming the assumption that Chinese users try to avoid expressing extreme emotions. Note that although Weibo has censorships, it does not include profanity filters.
@@ 6 Limitations
Translation: Lexica need to be adapted to the cultures to measure psychological phenomenon accurately. We tried using Chinese valence-arousal words (CVAW, Lee et al., 2022). However, we did not proceed further as the methods of building NRC-VAD (for English) and CVAW (for Chinese) lexica were different and could cause misalignment. We wanted to control for such differences by choosing a lexicon that has sufficient coverage while also being used in multiple prior works across both languages Wenjia et al. (2023); Mohammad (2016); Li et al. (2019); Chen et al. (2019); Das and Dutta (2021). Further, with over 20K entries, NRC-VAD is the largest manually created emotion lexicon that has translations across several languages.
Censorship: We acknowledge that censorship on Sina Weibo is a challenge. Despite this issue, Weibo has been successfully used across multiple studies to understand different psychological outcomes (e.g. affect, stress, depression; Pan et al., 2021; Tang et al., 2019; Li et al., 2014)
@@ 7 Ethics
The University of Pennsylvania’s Institutional Review Board declared this project exempt (IRB protocol # 829811).
@@
Appendix A Details of Pearson $r$
In our paper, Pearson $r$ correlation is independently calculated for post level valence and arousal scores, which gives each n-gram and topic a valence and arousal score. The Pearson r correlation coefficient is calculated with the OLS regression. Since the Coefficient in OLS regression is:
$\beta=\frac{Cov(x,y)}{Var(x)}$ (5)
And Pearson Correlation Coefficient is:
$r=\frac{Cov(x,y)}{\sqrt{Var(x)\cdot Var(y)}}$ (6)
So $\beta$ can be represented by $r$ with:
$\beta=r\cdot\frac{SD(y)}{SD(x)}$ (7)
So when normalized, $\beta=r$.
Figure 6: Example of Russell’s two-dimensional circumplex model. Captured from Russell 1980.
Appendix B Explanation of Conditional $R^{2}$
@@
The method we used for finding functional relationships between valence and arousal is a follow-up analysis using the same method used in the previous work. Table 2 in Kuppens et al., 2013 listed BIC, PostP, and the best $R^{2}$ for different studies and datasets, which are consistent with our result - relatively low $R^{2}$ scores. In our paper, we expand this work using social media data to see if a similar conclusion can hold across cultures with the functional relationship between valence and arousal.
@@
Appendix C Details of Valence-Arousal Relation Models
@@
Appendix D Removal of NRC-VAD Lexicons
@@
Appendix E Pre-Translated Weibo Figures